# Retrieval-Augmented Generation with Estimation of Source Reliability

### v4

@@ Abstract
Jeongyeon Hwang1, Junyoung Park 1, Hyejin Park1, Dongwoo Kim1, Sangdon Park1, Junseul Ok1,thanks: Corresponding authors. Email: jungseul@postech.ac.kr, 1Pohang University of Science and Technology (POSTECH), South Korea
@@ 2 Related Works
Robust RAG against misinformation. In response to misinformation risks in RAG, several robust methods have been proposed, primarily focusing on improving answer generation after retrieval. Weller et al. (2024); Xiang et al. (2024) utilize majority voting, which is effective only when most retrieved documents are trustworthy. Deng et al. (2024) evaluates document credibility using LLMs’ internal knowledge, but this approach is inherently limited as it misaligns with RAG’s core rationale of leveraging external knowledge to address LLMs’ limitations. Pan et al. (2024) assigns binary credibility scores (high/low) to retrieved documents based on source reputation and incorporates them into prompts. However, this approach is unsuitable for sources with obscure reputations, and reputation does not necessarily reflect actual reliability. In contrast, RA-RAG integrates source reliability into the retrieval and answer generation process by explicitly estimating it. Additional discussion on related works is presented in Appendix D.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that explicitly distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents originating from source $i\in[N]$. The definition of a “source” is application-dependent and may vary in granularity: sources can be fine-grained (e.g., individual social media accounts or statements by specific individuals such as politicians) or coarse-grained (e.g., news websites). In Section 6, we demonstrate practical applications of this framework.
@@ 5 Experiments
Figure 2: Accuracy performance on NQ and TQA datasets. (a) Results with heterogeneous reliability via beta priors for varying sources (4 to 9) across the Llama3-8B-Instruct and GPT-4o-mini models. See Appendix F.1 on the HotpotQA dataset and Phi3-mini-Instruct model. (b) Results with adversarial setting via adversary-hammer prior for varying adversaries (1 to 7) with Llama3-8B-Instruct model, highlighting overall trends. Exact values, which may overlap significantly, are provided in Appendix A4 with HotpotQA results.
@@ 5.1 Experimental setups
For analytical simplicity, we set $r_{i}=0.6$ for both priors to focus on evaluating $p_{i}$. The details of the data generation and source construction processes are provided in Appendix H. Due to the computational and financial constraints, we use 1,600 queries per dataset, allocating 200 queries for reliability estimation and 1,400 queries for test evaluation.
@@
Models. For language models, we use Llama3-8B-Instruct (Dubey et al., 2024), Phi3-mini-Instruct (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024), and Llama2-7B (Touvron et al., 2023). As a retriever, we use Contriever (Izacard et al., 2022). Due to space limitations, the results for Llama2-7B with Self-RAG fine-tuned on Llama2-7B are provided in Appendix E.
@@
Figure 3: A qualitative example comparing the answers produces by MV and RA-RAG for a query from the NQ dataset. Additional examples are available in Appendix G.
@@ 5.2 Main results
Beta prior. We evaluate RA-RAG across varying numbers of sources to assess its effectiveness in heterogeneous source reliability. As shown in Figure 2(a), RA-RAG consistently outperforms baselines, with performance gains increasing as more sources are incorporated. These results demonstrate the robustness of our approach in aggregating information from multiple sources with varying reliability. Notably, by selecting a subset of reliable and relevant sources using $\kappa$-RRSS, RA-RAG achieves performance comparable to Oracle WMV while improving efficiency by relying on fewer sources. In contrast, Robust RAG struggles with varying source reliability, as its certification assumption does not hold, resulting in lower performance than MV. Additionally, RA-RAG significantly outperforms Self-RAG, as shown in the Appendix E. These results emphasize the importance of differentiating between sources to prevent retrieval results from being overwhelmed by misinformation.
@@
Adversary-hammer prior. To evaluate the robustness of RA-RAG in the worst-case scenario, we use the adversary-hammer prior with a total of 9 sources on the NQ dataset with Llama3-8B-Instruct, as shown in Figure 2(b). Our RA-RAG demonstrates significant robustness against adversaries, whereas Robust RAG and Vanilla RAG suffer severe performance degradation as the number of adversaries increases. Similarly, Self-RAG experiences significant performance degradation, as detailed in Appendix E. Notably, when the number of adversaries exceeds four, the performance of MV significantly degrades due to the dominance of misinformation, leading MV to select incorrect answers.
@@ 5.3 Ablation studies and analysis
Impact of $\kappa$ for $\kappa$-RRSS. We conduct an ablation study on $\kappa$ across three datasets using Llama3-8B-Instruct with 9 sources. As shown in Figure 4, RA-RAG achieves stable performance starting from $\kappa=4$, indicating that selecting a small subset of reliable and relevant sources can maintain performance while significantly reducing computational overhead. This trend is consistent across other datasets; refer to Appendix F.2.
@@
As shown in Table 2, incorporating $\kappa$-RRSS consistently enhances computational efficiency across all metrics, with the reduction rate increasing as the number of sources grows. For example, in terms of token consumption, $\kappa$-RRSS reduces the total tokens processed by 2.6% with 5 sources, 32.3% with 10 sources, and 99.1% with 1000 sources. These significant efficiency gains indicate that $\kappa$-RRSS reduces computational overhead while maintaining reliable performance. Additional wall-clock time comparisons with baseline methods are provided in Appendix K. Further comparisons of accuracy between w/ and w/o $\kappa$-RRSS across different number of sources and models can be found in Appendix F.3.
@@ 6 Real-world Application
Figure 4: Accuracy across different values of $\kappa$ on Llama3-8B-Instruct and NQ dataset. Results for other datasets are provided in Appendix F.2
@@ 6.2 Experimental results
Figure 5: Results of reliability estimation under augmented variation for User A. Additional results for Politician A and B are in Appendix L.
@@ 7 Conclusion
We show that our reliability estimation method effectively estimates the reliability of real-world sources, particularly for news-related claims with abundant fact-checking sources. However, it remains challenging to apply to specialized topics due to limited references for cross-verification. Exploring expert knowledge as an alternative could help address this limitation and presents a promising direction for future work. Additionally, since our framework operates in an offline setting, it requires periodic updates to capture changes in source reliability over time. While such updates can be performed efficiently, as demonstrated in Appendix K.2, there remains a risk that even reliable sources may suddenly disseminate large volumes of unreliable information, such as in cases of account hacking. Although such cases are uncommon, this threat underscores the need for more responsive systems. A promising direction for future work is the development of an online framework that continuously updates reliability estimates in real time, enabling adaptive responses to the evolving information landscape.
@@ Appendix A Details of Misalignment Filtration
where $\mathcal{E}$ is the Alignscore function, $\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))=\mathcal{E}(\hat{y}_{i}^{*},\mathcal{R}(q,\mathcal{S}_{i}))$, and $\tau$ is the threshold. In all experiments, we set $\tau=0.1$ following Lei et al. (2023), which identifies this threshold as optimal for a real-world dataset comprising CNN and Daily Mail articles. For further analysis, we also conduct an ablation study on $\tau$ in Section M.
@@ Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
Table A1: Answer type distribution (%) by retrieved document type in the filtering $f_{\text{align}}$ ablation study on Llama3-8B-Instruct model and TQA dataset. Additional results for other datasets and models are provided in Appendix M.
@@ Appendix C Analysis of Filtering
Appendix D Extended Related Works
Retrieval-augmented generation. RAG enhances LLM performance by retrieving information from external databases, reducing hallucinations, and improving access to up-to-date knowledge (Lewis et al., 2020; Guu et al., 2020). Since irrelevant documents are prevalent in retrieval results, many studies have focused on enhancing RAG’s robustness through advanced retrieval methods, such as adaptive retrieval (Asai et al., 2024; Jiang et al., 2023), reranking retrieved documents (Glass et al., 2022), and query reformulation (Wang et al., 2023; Ma et al., 2023). While these approaches improve the retrieval process, they still rely on relevance measures between queries and documents, leaving them vulnerable to misinformation. Zou et al. (2024).
Appendix E Experiments Results on Llama2-7B Model
@@
Appendix F Additional Experimental Results
F.1 Beta prior and adversary-hammer prior
@@
F.2 Ablation study of $\kappa$ for $\kappa$-RRSS
@@
F.3 Ablation study of $\kappa$-RRSS in RA-RAG
@@
Appendix G Qualitative Examples
@@
Appendix H Benchmark of Multi-source RAG
@@
The specific prompts used to generate the data are provided in Appendix I.
@@
Appendix I Prompts for Constructing Multi-Source Benchmark
I.1 Prompt for factual data generation
@@
I.2 Prompt for misinformation generation
@@
Appendix J Instruction for Answer Generation
@@
Appendix K Computational efficiency with wall-clock time
We further evaluate computational efficiency by measuring wall-clock time for both inference (Section K.1) and source reliability estimation (Section K.2). Experiments were conducted using the Beta prior on the NQ dataset, as described in Section 5.1, with a single RTX 6000 Ada GPU.
K.1 Inference phase
@@
K.2 Reliablity estimation phase
@@
Appendix L Experimental Results on Estimating the Reliability of Real-World Sources
@@
Appendix M Extended ablation studies for filtering

### v5

@@ Abstract
Jeongyeon Hwang1, Junyoung Park 1, Hyejin Park1, Dongwoo Kim1, Sangdon Park1, Jungseul Ok1,thanks: Corresponding authors. Email: jungseul@postech.ac.kr, 1Pohang University of Science and Technology (POSTECH), South Korea
@@ 2 Related Works
Retrieval-augmented generation. Since irrelevant documents are prevalent in retrieval results, many studies have focused on enhancing RAG’s robustness through advanced retrieval methods, such as adaptive retrieval (Asai et al., 2024; Jiang et al., 2023), reranking retrieved documents (Glass et al., 2022), and query reformulation (Wang et al., 2023; Ma et al., 2023). While these approaches improve the retrieval process, they still rely on relevance measures between queries and documents, leaving them vulnerable to misinformation. Zou et al. (2024).
Robust RAG against misinformation. In response to misinformation risks in RAG, several robust methods have been proposed, primarily focusing on improving answer generation after retrieval. Weller et al. (2024); Xiang et al. (2024) utilize majority voting, which is effective only when most retrieved documents are trustworthy. Deng et al. (2024) evaluates document credibility using LLMs’ internal knowledge, but this approach is inherently limited as it misaligns with RAG’s core rationale of leveraging external knowledge to address LLMs’ limitations. Pan et al. (2024) assigns binary credibility scores (high/low) to retrieved documents based on source reputation and incorporates them into prompts. However, this approach is unsuitable for sources with obscure reputations, and reputation does not necessarily reflect actual reliability. In contrast, RA-RAG explicitly estimates source reliability and incorporates it into RAG systems.
Learning from noise sources. Learning from noisy sources has been extensively studied due to the scarcity of clean datasets in real-world applications Liu et al. (2012); Li and Yu (2014); Ok et al. (2016); Khetan et al. (2017); Zeng et al. (2018); Ok et al. (2019); Kim et al. (2022). A common approach is to estimate the reliability of data providers to aggregate trustworthy information from mixed-quality data. RAG systems face a similar challenge, as internet sources vary in reliability, but existing methods lack mechanisms for robust aggregation. To the best of our knowledge, this is the first work to explicitly embed reliability estimation to obtain robust information in RAG systems.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents from source $i\in[N]$. The definition of a “source” is application-dependent and may vary in granularity: sources can be fine-grained (e.g., individual social media accounts or statements by specific individuals such as politicians) or coarse-grained (e.g., news websites). In Section 6, we demonstrate practical applications of this framework.
@@ 5 Experiments
Figure 2: Accuracy performance on NQ and TQA datasets. (a) Results with heterogeneous reliability via beta priors for varying sources (4 to 9) across the Llama3-8B-Instruct and GPT-4o-mini models. See Appendix E.1 on the HotpotQA dataset and Phi3-mini-Instruct model. (b) Results with adversarial setting via adversary-hammer prior for varying adversaries (1 to 7) with Llama3-8B-Instruct model, highlighting overall trends. Exact values, which may overlap significantly, are provided in Appendix A4 with HotpotQA results.
@@ 5.1 Experimental setups
For analytical simplicity, we set $r_{i}=0.6$ for both priors to focus on evaluating $p_{i}$. The details of the data generation and source construction processes are provided in Appendix G. Due to the computational and financial constraints, we use 1,600 queries per dataset, allocating 200 queries for reliability estimation and 1,400 queries for test evaluation.
@@
Models. For language models, we use Llama3-8B-Instruct (Dubey et al., 2024), Phi3-mini-Instruct (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024), and Llama2-7B (Touvron et al., 2023). As a retriever, we use Contriever (Izacard et al., 2022). Due to space limitations, the results for Llama2-7B with Self-RAG fine-tuned on Llama2-7B are provided in Appendix D.
@@
Figure 3: A qualitative example comparing the answers produces by MV and RA-RAG for a query from the NQ dataset. Additional examples are available in Appendix F.
@@ 5.2 Main results
Beta prior. We evaluate RA-RAG across varying numbers of sources to assess its effectiveness in heterogeneous source reliability. As shown in Figure 2(a), RA-RAG consistently outperforms baselines, with performance gains increasing as more sources are incorporated. These results demonstrate the robustness of our approach in aggregating information from multiple sources with varying reliability. Notably, by selecting a subset of reliable and relevant sources using $\kappa$-RRSS, RA-RAG achieves performance comparable to Oracle WMV while improving efficiency by relying on fewer sources. In contrast, Robust RAG struggles with varying source reliability, as its certification assumption does not hold, resulting in lower performance than MV. Additionally, RA-RAG significantly outperforms Self-RAG, as shown in the Appendix D. These results emphasize the importance of differentiating between sources to prevent retrieval results from being overwhelmed by misinformation.
@@
Adversary-hammer prior. To evaluate the robustness of RA-RAG in the worst-case scenario, we use the adversary-hammer prior with a total of 9 sources on the NQ dataset with Llama3-8B-Instruct, as shown in Figure 2(b). Our RA-RAG demonstrates significant robustness against adversaries, whereas Robust RAG and Vanilla RAG suffer severe performance degradation as the number of adversaries increases. Similarly, Self-RAG experiences significant performance degradation, as detailed in Appendix D. Notably, when the number of adversaries exceeds four, the performance of MV significantly degrades due to the dominance of misinformation, leading MV to select incorrect answers.
@@ 5.3 Ablation studies and analysis
Impact of $\kappa$ for $\kappa$-RRSS. We conduct an ablation study on $\kappa$ across three datasets using Llama3-8B-Instruct with 9 sources. As shown in Figure 4, RA-RAG achieves stable performance starting from $\kappa=4$, indicating that selecting a small subset of reliable and relevant sources can maintain performance while significantly reducing computational overhead. This trend is consistent across other datasets; refer to Appendix E.2.
@@
As shown in Table 2, incorporating $\kappa$-RRSS consistently enhances computational efficiency across all metrics, with the reduction rate increasing as the number of sources grows. For example, in terms of token consumption, $\kappa$-RRSS reduces the total tokens processed by 2.6% with 5 sources, 32.3% with 10 sources, and 99.1% with 1000 sources. These significant efficiency gains indicate that $\kappa$-RRSS reduces computational overhead while maintaining reliable performance. Additional wall-clock time comparisons with baseline methods are provided in Appendix J. Further comparisons of accuracy between w/ and w/o $\kappa$-RRSS across different number of sources and models can be found in Appendix E.3.
@@
Figure 4: Accuracy across different values of $\kappa$ on Llama3-8B-Instruct and NQ dataset. Results for other datasets are provided in Appendix E.2
@@ 6.1 Setup
Figure 5: Results of reliability estimation under augmented variation for User A. Additional results for Politician A and B are in Appendix K.
6.2 Experimental results
@@ 7 Conclusion
We show that our reliability estimation method effectively estimates the reliability of real-world sources, particularly for news-related claims with abundant fact-checking sources. However, it remains challenging to apply to specialized topics due to limited references for cross-verification. Exploring expert knowledge as an alternative could help address this limitation and presents a promising direction for future work. Additionally, since our framework operates in an offline setting, it requires periodic updates to capture changes in source reliability over time. While such updates can be performed efficiently, as demonstrated in Appendix J.2, there remains a risk that even reliable sources may suddenly disseminate large volumes of unreliable information, such as in cases of account hacking. Although such cases are uncommon, this threat underscores the need for more responsive systems. A promising direction for future work is the development of an online framework that continuously updates reliability estimates in real time, enabling adaptive responses to the evolving information landscape.
@@ Appendix A Details of Misalignment Filtration
where $\mathcal{E}$ is the Alignscore function, $\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))=\mathcal{E}(\hat{y}_{i}^{*},\mathcal{R}(q,\mathcal{S}_{i}))$, and $\tau$ is the threshold. In all experiments, we set $\tau=0.1$ following Lei et al. (2023), which identifies this threshold as optimal for a real-world dataset comprising CNN and Daily Mail articles. For further analysis, we also conduct an ablation study on $\tau$ in Section L.
@@ Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
Table A1: Answer type distribution (%) by retrieved document type in the filtering $f_{\text{align}}$ ablation study on Llama3-8B-Instruct model and TQA dataset. Additional results for other datasets and models are provided in Appendix L.
@@ Appendix C Analysis of Filtering
Appendix D Experiments Results on Llama2-7B Model
@@
Appendix E Additional Experimental Results
E.1 Beta prior and adversary-hammer prior
@@
E.2 Ablation study of $\kappa$ for $\kappa$-RRSS
@@
E.3 Ablation study of $\kappa$-RRSS in RA-RAG
@@
Appendix F Qualitative Examples
@@
Appendix G Benchmark of Multi-source RAG
@@
The specific prompts used to generate the data are provided in Appendix H.
@@
Appendix H Prompts for Constructing Multi-Source Benchmark
H.1 Prompt for factual data generation
@@
H.2 Prompt for misinformation generation
@@
Appendix I Instruction for Answer Generation
@@
Appendix J Computational efficiency with wall-clock time
We further evaluate computational efficiency by measuring wall-clock time for both inference (Section J.1) and source reliability estimation (Section J.2). Experiments were conducted using the Beta prior on the NQ dataset, as described in Section 5.1, with a single RTX 6000 Ada GPU.
J.1 Inference phase
@@
J.2 Reliablity estimation phase
@@
Appendix K Experimental Results on Estimating the Reliability of Real-World Sources
@@
Appendix L Extended ablation studies for filtering